OpenStudy (anonymous):

-2(x + 3) = -2(x + 1) - 4.

OpenStudy (anonymous):

i think it is no solution

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Hey get photo math on your device, and can I get a medal because that will help you with a lot of questions

OpenStudy (anonymous):

And it will give you steps they took to do it

OpenStudy (anonymous):

is it free

OpenStudy (anonymous):

Pretty sure it still is

OpenStudy (anonymous):

can you just check it cause i know how to do it but im confused cause it cancels out

OpenStudy (anonymous):

I'm pretty sure there is no solution to this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

One sec tho

OpenStudy (anonymous):

Yea I'm pretty sure there is no solution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is wrong it is like a+b=b+a

OpenStudy (anonymous):

That would end up being any number -2(x + 3) = -2(x + 1) - 4. That would turn into -2x+-6=-2x+(-2)-4. So you would add the -2 and -4. Then you would get -2x+(-6)= -2x+(-6)